OpenStudy (anonymous):

Given that f(x)=X2+3x+6 and g(x) = x-3/2 , solve for f(g(x)) when x=1 What's the answer ?

OpenStudy (dan815):

f(g(x)) <--- replace all the x in f(x) with g(x)

OpenStudy (dan815):

when x=1, f(g(1)) = f(1-3/2)=f(-1/2)=(-1/2)^2+3(-1/2)+6=-5/4+6=19/4

OpenStudy (anonymous):

So the answer is 4 ?

OpenStudy (mathmale):

Kevin, it's tempting to zero in on the answer, but you'd benefit much more by actually doing and sharing the work necessary to calculate that answer. What are your thoughts? What have you done so far in working on this problem? What does f(g(x)) mean to you?